Grade 12Physical Chemistry

What is the concentration of O2 in a freshwater stream in equilibrium with air at 25oC and 1.0 bar?
Express the answer in grams of O2 per kg of solvent. (Henry’s law constant for O2 is 1.3x10-3 mol/kgxbar

Shakti Nirmal
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
The mole fraction of O2 in air is 0.21 and assuming the total pressure is 1.0bar, the partial pressure of O2 is 0.21bar. Using this pressure for Pg in Henry`s law we have...Solubility of O2 = KHPg = 1.3×10−3mol/Kg.bar (0.21 bar) = 2.7 × 10-4 mol/lThis concentration in grams per liter can then be calculated using the molar mass of O2.Solubility of O2 = 2.7×10−4mol/Kg × 32.0g/mol = 0.0087 g/KgThis concentration of O2 (8.7 mg/Kg) is quite low, but it is sufficient to provide the oxygen required by aquatic life.